Application of Overhead

Budgeted Total Manufacturing Overhead for 2020 =$400,000

Budgeted Total Machine hours for 2020=50,000

POHR= ?

Actual Use of Machine Hours for Two jobs were as follows:

Job #1 machine hours-1600 hours

Job #2 machine hours-1000 hours

Total machine hours=?

What is the Manufacturing Overhead Applied for each job?

What is the Total Manufacturing Overhead Applied?

Show ALL work
Answer #1

Budgeted Total Manufacturing Overhead for 2020 = $400,000

Budgeted Total Machine hours for 2020 = 50,000

Predetermined overhead rate = Budgeted Total Manufacturing Overhead for 2020/ Budgeted Total Machine hours for 2020

= 400,000/50,000

= $8 per machine hour

Actual machine hours used for Job#1 = 1,600 hours

Actual machine hours used for Job#2 = 1,000 hours

Total machine hours = Actual machine hours used for Job#1 + Actual machine hours used for Job#2

= 1,600+1,000

= 2,600

Manufacturing overhead applied for Job#1 = Actual machine hours used for Job#1 x Predetermined overhead rate

= 1,600 x 8

= $12,800

Manufacturing overhead applied for Job#2 = Actual machine hours used for Job#2 x Predetermined overhead rate

= 1,000 x 8

= $8,000

Total Manufacturing Overhead Applied = Manufacturing overhead applied for Job#1+ Manufacturing overhead applied for Job#2

= 12,800+8,000

= $20,800

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.Thanks!!!